OpenStudy (anonymous):

Which expression could be a calculator’s display for the number 0.00000000000096? A. 9.6e13 B. 96e12 C. 9.6e−13 D. 0.96e−14

OpenStudy (anonymous):

@shinalcantara

OpenStudy (shinalcantara):

just count the number of zeroes to the right until the decimal point is in between 9 & 6.. since you're converting it to a mathematical number then the exponent would be negative

OpenStudy (anonymous):

ok i got 2 more

OpenStudy (anonymous):

What is 7.8 × 105 m/h in kilometers per day? A. 3.24 × 1010 km/day B. 3.24 × 108 km/day C. 1.872 × 106 km/day D. 1.872 × 104 km/day

OpenStudy (shinalcantara):

REMEMBER -that when you convert a certain number to a mathematical number, moving the decimal place to the right makes the exponent negative while moving it to the left makes a positive exponent. -that when you convert a mathematical number to it's standard form, a positive exponent implies that you have to move the decimal place to the right while moving it to the left when it's exponent is negative ONE MORE THING -the coefficient to consider when expressing the number into mathematical notation must be equal to or greater than one but less than 10.. :)

OpenStudy (shinalcantara):

what's your answer for the question above btw?

OpenStudy (anonymous):

i think its b hold on which one

OpenStudy (shinalcantara):

the question involving mathematical notation

OpenStudy (anonymous):

i think its d

OpenStudy (shinalcantara):

it's actually c..

OpenStudy (anonymous):

what grade you in

OpenStudy (anonymous):

What is 3.2 × 10−4 m/h in centimeters per second? A. 8.889 × 10−6 cm/s B. 8.889 × 10−5 cm/s C. 1.152 × 10−5 cm/s D. 1.152 × 10−7 cm/s

OpenStudy (shinalcantara):

i'm actually a bachelor's degree holder

OpenStudy (anonymous):

oh

OpenStudy (shinalcantara):

try to recheck for the given that '7.8 x 105 m/hr' the answers are quite not appropriate

OpenStudy (shinalcantara):

ah i see.. the given is \[7.8 \times 10 ^{5}\frac{ m }{ hr } (\frac{ 24 hrs }{ 1 day })(\frac{ 1km }{ 1000m }) = ?\]

OpenStudy (shinalcantara):

\[3.2 \times 10^{-4}\frac{ m }{ hr }(\frac{ 100 cm }{ 1 m })(\frac{ 1 hr }{ 3600s }) = ?\]

OpenStudy (anonymous):

for wat question